I am having a hard time figuring out this problem can you help me understand it better.
I need to figure out the surface area of the solid right circular cylinder with
r = 9cm and h = 7cm
Would my surface area be 904.3cm^2
Thanks :o)

Answer by stanbon(75887)

figure out the surface area of the solid right circular cylinder with
r = 9cm and h = 7cm
Would my surface area be 904.3cm^2
----
Surface area = side + 2*ends
side area = perimeter * height = 2(pi)9*7 = 126pi cm^2
2 ends = 2[pi*9^2] = 162pi cm^2
----
Total surface area = 288pi cm^2 is approximately 904.78 cm^2
========
Cheers,
Stan H.
===========